x=log bc base a, y=log ca base b,z=log ab base c prove that x+y+z=xyz-2
1
Expert's answer
2019-09-02T09:26:33-0400

Let


"x=\log_a(bc), \; y=\log_b(ac),\: z=\log_c(ab)"

We put

"\log_a(b)=u,\:\log_a(c)=v, \; \log_b(c)=w"

Thus


"x*y*z=[u+v]*[1\/u+w]*[1\/v+1\/w]"

"=[1+v\/u+uw+vw]*[1\/v+1\/w]"

"=1\/v+1\/u+uw\/v+w+1\/w+v\/(uw)+u+v"

Since


"\log_a(b)*\log_b(c)=\log_a(c)"

we get


"v\/(uw)=1"

So


"x*y*z=1\/v+v+1\/u+u+1\/w+w+2"

Also we have


"x+y+z=[u+v]+[1\/u+w]+[1\/v+1\/w]=1\/v+v+1\/u+u+1\/w+w"

Finally

"x+y+z=x*y*z-2"